### v4

@@ 4.4 Complexity Analysis
This section examines the time complexity associated with each category of circuits. For linear circuits, including those characterized by $K$-coefficient linear circuits, the Gauss-Jordan elimination method has a time complexity of $O(n^{3})$ [3]. Our statistical analysis of Circomlib [32] indicates that the majority of cases involve linear circuits. Consequently, we anticipate that our algebraic-based method will significantly enhance performance compared to SMT solvers. In the case of higher-order circuits, we employed the $F_{5}$ algorithm, which is designed for computing Gröbner bases.[23] This algorithm typically exhibits $O(2^{n})$ or $O(2^{2^{n}})$ time complexity in the worst-case scenarios. However, the $F_{5}$ algorithm demonstrates substantial improvements over its predecessors, such as Buchberger’s algorithm[18] and the $F_{4}$ algorithm.
@@ 5.1 Environment Setting and Metrics
Environment configuration. The evaluations were performed on a machine with a 16-core AMD Ryzen 9 5950x processor and 128GB of DRAM.
Check Benchmark. We settle a representative benchmark to evaluate the effectiveness and accuracy of $\text{AC}^{4}$ and Picus [46]. The benchmark was based on the data set provided by $\text{AC}^{4}$ and Picus, which had been collected from Circomlib, the standard library from Circom. All cases of our benchmark were divided into two parts: Circomlib-utils benchmarks, which was extracted from the Circomlib utils library and had a smaller size, and Circomlib-core benchmarks, which was extracted from the Circomlib core library and had a larger size [46].
Furthermore, we established a benchmark to evaluate the effectiveness and accuracy of AC4 and the halo2-analyzer[31], as implemented in the study by Soureshjani et al. (2023) [52]. This benchmark was constructed using a dataset of classical halo2 circuits, which were collected from various GitHub projects.
@@
Benchmark Metrics Checked Ratio. The solved benchmarks are initially categorized into three types: precisely solved, algebraically solved, and unknown. The precisely solved cases are further categorized into three types: precisely exact-constrained, which means that the circuit is safe; precisely overconstrained, indicating that the circuit is overconstrained; and precisely underconstrained, which suggests that the circuit is underconstrained. We omit this metric as $PS$ (Precisely Solved).
@@ 5.2 Detailed Evaluation over Circom Circuits
We categorize circuits based on their constraint numbers in Table 2, considering those with a constraint number less than 100 as “small” circuits, those with a constraint number in the range $[100,1000)$ as “medium” circuits, and those with a constraint number in the range $[1000,\infty)$ as “large” circuits. This classification, which may be rudimentary, does not consider output signals. It is evident that small circuits can be solved directly, whereas other circuits may require specific restrictions to be solved within an acceptable time frame.
@@
Checking time. The checking time of the benchmark set is presented in Table 3, with a focus on the relationship between benchmark size and solving difficulty. It is observed that the larger benchmarks present greater difficulty in solving, resulting in longer solution times. However, in some large circuits, despite a high number of constraints, the number of output signals is limited to 1, allowing for a relatively easy check once the file read operation concludes. Nonetheless, there exist extremely large circuits that are characterized by a linear type, yet their extensive constraint size renders the Gaussian-Jordan method challenging to execute, leading to out-of-memory issues. Conversely, smaller cases are relatively straightforward to solve directly but due to limitations, the precise solution is not attainable, resulting in algebraic outcomes.
@@
Table 3: The efficiency benchmark of $AC^{4}$ is categorized by size and includes the time for precisely solved and algebraically solved results. The terms K, U, O, v, PS, AS represent known, unknown, output, variables, precisely solved, and algebraically solved, respectively. The explanation of PS, AS is discussed in Section 5.1.
@@
Avg. PS Time(s) 0.57 0.52 9.06 1.83 0.42 2.4 14.59 3.71 3.03
Avg. AS&PS Time(s) 1.26 2.45 8.33 2.41 0.40 2.4 13.47 3.48 3.06
Checked Ratio. The checked ratio of the benchmark sets categorized by circuit type is presented in Table 4. The data indicates a significantly high solved ratio for precisely linear circuits. However, Substantial constraints in circuits result in very large augmented matrices. As a consequence, the solve ratio of $k-$coefficient circuits is relatively low. This is attributed to the adoption of an algebraic strategy for $k-$coefficient circuits, which in turn leads to the direct solving of small-size $k-$coefficient circuits without returning a precise result.
@@
Table 4: Main Experiment results, we categorize circuits based on their types: pl means precisely linear type, kcl means k coefficient linear, q means quadratic, ps means precisely solved, as means algebraically solved.
@@
Comparison with Baseline
In the paper [46], Picus, an implementation of the tool $\textbf{QED}^{2}$, is utilized for verifying the uniqueness property (under-constrained signals) of ZKP circuits. This implementation employs static program analysis and SMT solvers checking to achieve the verification of the uniqueness of ZKP circuits. Subsequently, we benchmarked the Picus program using the same benchmark set.
The completed Picus ratio for the utils benchmark set is 80.33%, for the core benchmark set it is 62.61%, and for both sets it is 69.36%. Additionally, the average solving time for the Picus solver is 425.73 seconds for the utils benchmark set, 317.88 seconds for the core benchmark set, and 365.52 seconds for both sets.
(a) Comparison between the checking time of $\text{AC}^{4}$ and Picus.
(b) Comparison between checked ratio of $\text{AC}^{4}$ and Picus
Figure 4: The checking time and checked ratio comparisons between $\text{AC}^{4}$ and Picus
The comparison of the two tools based on the solved ratio can be found in Fig. 4b. Our advantage lies in precisely linear cases as well as some heuristic methods for $k-$coefficient and binary cases.Fig. 4b shows the solving time comparison of the two tools.
@@ 5.3 Evaluation over Halo2 Circuits
Given the current dearth of comprehensive research on halo2 circuits and the absence of template circuits provided by the official Zcash organization, we undertook the task of extracting 10 of the most frequently utilized halo2 circuit types from GitHub repositories dedicated to halo2 projects. These extractions were subsequently employed to construct a relevant dataset. Utilizing this custom-curated dataset, we conducted benchmark analyses in conjunction with the halo2-analyzer.
As shown in figure 5, AC4 demonstrates a enhancement in efficiency and accuracy compared to halo2-analyzer, exhibiting a 26.73% improvement in the verification time and 10% improvement in the checked ratio for halo2 circuits.
@@
(b) Comparison between checked ratio of $\text{AC}^{4}$ and Halo2-analyzer
Figure 5: The checking time and checked ratio comparisons between $\text{AC}^{4}$ and halo2-analyzer
@@ 5.4 Evaluation Result Analysis
Table 3 illustrates a significant correlation between the size of the circuit and the time required for checking. Despite their size, large circuits keep checking times low due to two critical factors. Firstly, the predominance of linear circuits, efficiently addressed by the Gauss-Jordan solver, contributes to the efficiency in solving a majority of cases. Additionally, optimizations targeting binary cases within k-coefficient and higher-order circuits serve to further reduce the solving time despite their slightly larger size.
The checked ratio is high for precisely linear cases; however, there are some cases for which we cannot provide precisely checked results due to the lack of an algebraic solution. Special input scenarios can reduce the augmented matrix’s rank, yielding a solution. In contrast, for the $K$-coefficient cases, the checked ratio is relatively low, even when a mixed solver is employed. This is because we did not implement certain static analysis methods to elaborate the benchmark set, which would have facilitated the checking of circuits. Conversely, the checked ratio is relatively high for quadratic cases, as most of these cases are in binary form. The detection method and optimization for binary quadratic circuits have significantly enhanced efficiency in checking these cases.
In the context of quadratic cases, Picus and halo2-analyzer employ the cvc5-ff-range solver. This solver first implements finite field theory and then utilizes Buchberger’s algorithm along with triangular decomposition to simplify polynomial equations. To address the problem, it further applies Collins’ cylindrical algebraic decomposition (CAD) [22] [43]. It is important to note that the time complexity of CAD is $O(2^{2^{n}})$, which can be prohibitive in many scenarios involving large quadratic cases. Additionally, the $F_{5}$ algorithm demonstrates superior performance over modular integers in various cryptographic challenges.
@@
Failure Analysis. Picus utilized their algorithm Uniqueness Constraint Propagation to solve all the $k-$coefficient circuits. The algorithm simplifies verification/resolution by strategically assigning values to unknown variables through partial equation analysis, reducing unknowns. However, our methods forgo static analysis, limiting $K$-coefficient case checking.
Analysis over halo2 circuits Halo2-analyzer employs cvc5-ff to verify circuit uniqueness, giving our tool an advantage in solving large-scale polynomial equation systems.
@@ 6 Related Work
Verification on ZK programs. Circomspect [14] serves as a static analyzer and linter designed for the Circom programming language. Picus [46] employs a static analysis method combined with an SMT solver. Circomspect and Picus share certain similarities with our work. However, Circomspect can only annotate a portion of underconstrained variables.
ZK circuit formalization has seen notable advancements: Shi et al. [50] proposed a data-flow-based algorithm for R1CS standardization, to enhance performance across R1CS tasks. Wen et al. [57] formalized and classified Circom vulnerabilities, introducing a CDG-based static analysis framework. CODA [37] emerged as a statically typed language for formal ZK system property specification and verification. Soureshjani et al.[52] applied abstract interpretation to examine halo2 circuit properties, on the Plonkish matrix using an SMT solver.
Research on ZK compilers with theorem proving is ongoing. Bangerter and Almeida et al.[7, 6, 2, 8], pioneered verifiable compilation using Isabelle/HOL [42]. PinocchioQ [25] verified Pinocchio[47], a C-based zk-SNARK, with Compcert [36] based on Coq [13]. Chin and Coglio et al [20, 21]. employed ACL2 [34] on Leo, while Bailey et al. [5] established formal proofs for six zk-SNARK constructions using the Lean theorem prover [41].
Algebra computation in circuit verification. Algebraic computation is a long-standing but niche endeavor in verification. Watanabe [56] utilized Gröbner Basis [24, 23] and polynomial reduction techniques to verify arithmetic circuits in computing and signal processing systems. Similarly, Lv et al. [40, 39] employed a computer-algebra-based approach for the formal verification of hardware. Scholl et al. [49] advanced this field by combining Computer Algebra with a Boolean SAT solver to verify divider circuits. In the context of Zero Knowledge Proof (ZKP) compilers, Alex et al. [45] presented a partial verification of a field-blasting compiler pass. Furthermore, Thomas et al. [30] introduced a novel automated reasoning method for determining the satisfiability of non-linear equation systems over finite fields. Additionally, Thomas et al. [29] enhanced the Yices2 SMT solver by implementing a new MCSat-based reasoning engine for finite fields.
Alex et al. [43] employed the C++ CLN library [17] to implement finite fields and utilized COCOALib’s [1] implementation of the Buchberger algorithm for preprocessing equations, subsequently applying CAD algorithm[22] to solve them and presents a new SMT solver for finite field equations that improves efficiency on cryptosystem verification tasks by using multiple simplified Gröbner bases and specialized propagation algorithms[44].

### v5

@@ 4.4 Complexity Analysis
This section examines the time complexity associated with each category of circuits. For linear circuits, including those characterized by $K$-coefficient linear circuits, the Gauss-Jordan elimination method has a time complexity of $O(n^{3})$ [ANDREN2007428]. Our statistical analysis of Circomlib [Iden3CircomlibLibrary] indicates that the majority of cases involve linear circuits. Consequently, we anticipate that our algebraic-based method will significantly enhance performance compared to SMT solvers. In the case of higher-order circuits, we employed the $F_{5}$ algorithm, which is designed for computing Gröbner bases.[f5algorithm] This algorithm typically exhibits $O(2^{n})$ or $O(2^{2^{n}})$ time complexity in the worst-case scenarios. However, the $F_{5}$ algorithm demonstrates substantial improvements over its predecessors, such as Buchberger’s algorithm[BUCHBERGER2006475] and the $F_{4}$ algorithm.
@@ 5.1 Environment Setting and Metrics
Environment configuration. The evaluations were performed on a machine with a 12th Gen Intel(R) Core(TM) i5-12500H processor and 16GB of DRAM.
Check Benchmark. We settle a representative benchmark to evaluate the effectiveness and accuracy of $\text{AC}^{4}$ ,Picus [10.1145/3591282] and CIVER [CIVER]. The benchmark was based on the data set collected from Circomlib, the standard library from Circom. All cases of our benchmark were divided into two parts: Circomlib-utils benchmarks, which was extracted from the Circomlib utils library and had a smaller size, and Circomlib-core benchmarks, which was extracted from the Circomlib core library and had a larger size [10.1145/3591282].
Furthermore, we established a benchmark to evaluate the effectiveness and accuracy of AC4 and the halo2-analyzer[halo2analyzer], as implemented in the study by Soureshjani et al. (2023) [SoureshjaniHJKG23]. This benchmark was constructed using a dataset of classical halo2 circuits, which were collected from various GitHub projects.
@@
Benchmark Metrics solved rate. The solved benchmarks are initially categorized into three types: precisely solved, algebraically solved, and unknown. The precisely solved cases are further categorized into three types: precisely exact-constrained, which means that the circuit is safe; precisely overconstrained, indicating that the circuit is overconstrained; and precisely underconstrained, which suggests that the circuit is underconstrained. We omit this metric as $PS$ (Precisely Solved).
@@ 5.2 Detailed Evaluation over Circom Circuits
We categorize circuits based on their constraint numbers in Table 2, considering those with a constraint number less than 100 as “small” circuits, those with a constraint number in the range $[100,1000)$ as “medium” circuits, and those with a constraint number in the range $1000,\infty)$ as “large” circuits. This classification, which may be rudimentary, does not consider output signals. It is evident that small circuits can be solved directly, whereas other circuits may require specific restrictions to be solved within an acceptable time frame.
@@
Results based on Circuit Size. Table 3 presents the results across circuits of different sizes, highlighting the relationship between benchmark scale and solving difficulty. In general, larger benchmarks exhibit greater solving difficulty, leading to longer solution times and lower checking rates.However, in the circomlib-utils benchmark, the “small” category exhibits a higher average solving time. This is because two small circuits, BitSub@binsub and BitSum@binsum, which are quadratic, take unusually long to solve (128.26 s and 185.88 s, respectively). If we exclude these two circuits, the average PS time and the average PS&AS time drop to 0.04 s and 0.89 s, respectively.
In some large circuits, despite a high number of constraints, the number of output signals is limited to 1, allowing for a relatively easy check once the file read operation concludes. Nonetheless, there exist extremely large circuits that are characterized by a linear type, yet their extensive constraint size renders the Gaussian-Jordan method challenging to execute, leading to out-of-memory issues. Conversely, smaller cases are relatively straightforward to solve directly but due to limitations, the precise solution is not attainable, resulting in algebraic outcomes.
@@
Table 3: Results of $AC^{4}$ categorized by circuit size, including the checking time and solved rate for precisely solved and algebraically solved cases. The terms K, U, O, v, PS, AS represent known, unknown, output, variables, precisely solved, and algebraically solved, respectively. The explanation of PS, AS is discussed in Section 5.1.
@@
Avg. PS Time(s) 7.89 0.20 7.03 6.79 0.10 1.36 10.94 2.55 1.08
Avg. AS&PS Time(s) 7.27 1.69 6.40 6.47 0.09 1.36 10.06 2.38 1.72
ps_rat 0.73 0.88 0.89 0.77 0.83 1.00 0.91 0.88 0.84
ps&as_rat 0.92 1.00 1.00 0.94 0.92 1.00 1.00 0.95 0.97
Result based on Circuit Type. Table 4 presents the results across circuits of different types, the data indicates a significantly high solved rate for precisely linear circuits. However, Substantial constraints in circuits result in very large augmented matrices. As a consequence, the checke rate of $k$-coefficient circuits is relatively low. This is attributed to the adoption of an algebraic strategy for $k$-coefficient circuits, which in turn leads to the direct solving of small-size $k$-coefficient circuits without returning a precise result.
@@
Table 4: Results of $AC^{4}$ categorized by circuit type, including the checking time and solved rate for precisely solved and algebraically solved cases. We categorize circuits based on their types: pl means precisely linear type, kcl means k coefficient linear, q means quadratic, ps means precisely solved, as means algebraically solved.
@@
Avg. PS Time(s) 1.30 0.05 44.93 6.79 3.00 0.03 1.39 2.55 2.05
Avg. AS&PS Time(s) 1.50 0.43 29.37 6.47 2.96 0.04 1.19 2.38 2.16
Comparison with Picus and CIVER
We compared our tool, $AC^{4}$, with Picus and CIVER using the circomlib-utils and circomlib-core benchmarks. For Picus, the average checking times for small, medium, and large circuits are 1.95 seconds, 6.32 seconds, and 3256.48 seconds, respectively, with corresponding solving rates of 80.24%, 48.32%, and 68.55%. The overall average checking time and solving rate are 365.52 seconds and 69.36%, respectively. For CIVER, the average checking times for small, medium, and large circuits are 0.85 seconds, 2.58 seconds, and 6.79 seconds, respectively, while the corresponding solving rates are 61.32%, 46.66%, and 58.33%. The overall average checking time and solving rate for CIVER are 1.19 seconds and 58.13%, respectively.
(a) Comparison between the checking time of $\text{AC}^{4}$, Picus and Civer.
(b) Comparison between solved rate of $\text{AC}^{4}$, Picus and Civer.
Figure 4: The checking time and solved rate comparisons between $\text{AC}^{4}$, Picus and Civer
As a comparison with $AC^{4}$, the checking time results are shown in Fig. 4a, and the checking rate results are shown in Fig. 4b. The results demonstrate that our tool achieves the highest checking rate among the three tools, particularly on medium- and large-sized circuits, indicating its superior capability in solving complex circuits. However, $AC^{4}$ exhibits a relative disadvantage in solving time for small and large circuits compared with CIVER. This is primarily because CIVER imposes a strict timeout limit, which causes it to skip overly complex circuits. These timeouts occur more frequently in large size circuits, leading to shorter average solving times but a lower overall checking rate. For the small circuits, the higher average time is largely due to two circuits, BitSub@binsub and BitSum@binsum. If we exclude these two circuits, the average time drops to 0.42 s, which is faster than CIVER.
@@ 5.3 Evaluation over Halo2 Circuits
In the absence of a standard library akin to Circomlib for Halo2, we meticulously selected 29 representative circuit implementations from several prominent open-source GitHub projects. These implementations encompass essential application scenarios, including large number arithmetic, multiplication and inner product, the Poseidon hash algorithm, and Ethereum Virtual Machine operations. The dataset is specifically categorized as follows: 10 test cases for small-scale circuits, 11 test cases for medium-scale circuits, and 8 test cases for large-scale circuits.
This curated dataset was then utilized to construct a comprehensive dataset for analysis. Using this dataset, we performed benchmark analyses in conjunction with the halo2-analyzer.
As shown in figure 5, AC4 demonstrates a enhancement in efficiency and accuracy compared to halo2-analyzer, exhibiting an improvement over efficiency and accuracy.
@@
(b) Comparison between solved rate of $\text{AC}^{4}$ and Halo2-analyzer
Figure 5: The checking time and solved rate comparisons between $\text{AC}^{4}$ and halo2-analyzer
@@ 5.4 Evaluation Result Analysis
Table 3 illustrates a significant correlation between the size of the circuit and the time required for checking and solved rate. Despite their size, large circuits keep checking times low due to two critical factors. Firstly, the predominance of linear circuits, efficiently addressed by the Gauss-Jordan solver, contributes to the efficiency in solving a majority of cases. Additionally, optimizations targeting binary cases within k-coefficient and higher-order circuits serve to further reduce the solving time despite their slightly larger size.
The solved rate is high for precisely linear cases; however, there are some cases for which we cannot provide precisely checked results due to the lack of an algebraic solution. Special input scenarios can reduce the augmented matrix’s rank, yielding a solution. In contrast, for the $K$-coefficient cases, the solved rate is relatively low, even when a mixed solver is employed. This is because we did not implement certain static analysis methods to elaborate the benchmark set, which would have facilitated the checking of circuits. Conversely, the solved rate is relatively high for quadratic cases, as most of these cases are in binary form. The detection method and optimization for binary quadratic circuits have significantly enhanced efficiency in checking these cases.
A more intriguing observation is that Table 4 checking time for the precisely linear type is unexpectedly large. This occurs because the distribution of circuit types in the benchmark is highly unbalanced—over 71.6% of the circuits are classified as precisely linear. Consequently, this category contains a larger proportion of large-sized circuits, which leads to the observed increase in average checking time.
In the context of quadratic cases, Picus and halo2-analyzer employ the cvc5-ff-range solver. This solver first implements finite field theory and then utilizes Buchberger’s algorithm along with triangular decomposition to simplify polynomial equations. To address the problem, it further applies Collins’ cylindrical algebraic decomposition (CAD) [Collins-CAD] [cvc5-ff]. It is important to note that the time complexity of CAD is $O(2^{2^{n}})$, which can be prohibitive in many scenarios involving large quadratic cases. Additionally, the $F_{5}$ algorithm demonstrates superior performance over modular integers in various cryptographic challenges.
@@
Limitations. Picus utilized their algorithm Uniqueness Constraint Propagation to solve all the $k$-coefficient circuits. The algorithm simplifies verification/resolution by strategically assigning values to unknown variables through partial equation analysis, thereby reducing the number of unknowns. In contrast, our method forgoes static analysis, which limits the checking of the $K$-coefficient cases.
Analysis over halo2 circuits. Halo2-analyzer employs cvc5-ff to verify circuit uniqueness, giving our tool an advantage in solving large-scale polynomial equation systems.
@@ 6 Related Work
Verification on ZK programs. Circomspect [Circomspect] serves as a static analyzer and linter designed for the Circom programming language. Picus [10.1145/3591282] employs a static analysis method combined with an SMT solver. Circomspect and Picus share certain similarities with our work. However, Circomspect can only annotate a portion of underconstrained variables.
ZK circuit formalization has seen notable advancements: Shi et al. [shi2023dataflowbased] proposed a data-flow-based algorithm for R1CS standardization, to enhance performance across R1CS tasks. Wen et al. [cryptoeprint:2023/190] formalized and classified Circom vulnerabilities, introducing a CDG-based static analysis framework. CODA [cryptoeprint:2023/547] emerged as a statically typed language for formal ZK system property specification and verification. Soureshjani et al.[SoureshjaniHJKG23] applied abstract interpretation to examine halo2 circuit properties, on the Plonkish matrix using an SMT solver.
Research on ZK compilers with theorem proving is ongoing. Bangerter and Almeida et al. [cryptoeprint:2008/471, 10.1007/978-3-642-16441-5_5, 10.1007/978-3-642-15497-3_10, bangerterdesign], pioneered verifiable compilation using Isabelle/HOL [ref1]. PinocchioQ [7536381] verified Pinocchio[10.1145/2856449], a C-based zk-SNARK, with Compcert [10.1145/1111037.1111042] based on Coq [Bertot2004]. Chin and Coglio et al [cryptoeprint:2023/1278, Coglio_2023]. employed ACL2 [588534] on Leo, while Bailey et al. [cryptoeprint:2023/656] established formal proofs for six zk-SNARK constructions using the Lean theorem prover [10.1007/978-3-319-21401-6_26].
Algebra computation in circuit verification. Algebraic computation is a long-standing but niche endeavor in verification. Watanabe [4601876] utilized Gröbner Basis [FAUGERE199961, f5algorithm] and polynomial reduction techniques to verify arithmetic circuits in computing and signal processing systems. Similarly, Lv et al. [6113989, 6167783] employed a computer-algebra-based approach for the formal verification of hardware. Scholl et al. [9218721] advanced this field by combining Computer Algebra with a Boolean SAT solver to verify divider circuits. In the context of Zero Knowledge Proof (ZKP) compilers, Alex et al. [DBLP:conf/cav/boundedFFB] presented a partial verification of a field-blasting compiler pass. Furthermore, Thomas et al. [DBLP:conf/lpar/HaderRK23] introduced a novel automated reasoning method for determining the satisfiability of non-linear equation systems over finite fields. Additionally, Thomas et al. [DBLP:conf/ijcar/HaderKIGK24] enhanced the Yices2 SMT solver by implementing a new MCSat-based reasoning engine for finite fields.
Alex et al. [cvc5-ff] employed the C++ CLN library [cln-cpp-lib] to implement finite fields and utilized COCOALib’s [COCOA-lib] implementation of the Buchberger algorithm for preprocessing equations, subsequently applying CAD algorithm[Collins-CAD] to solve them and presents a new SMT solver for finite field equations that improves efficiency on cryptosystem verification tasks by using multiple simplified Gröbner bases and specialized propagation algorithms[DBLP:conf/cav/SplitGrobnerBases].